OpenStudy (anonymous):

MIDTERM solution formulae for question 2

OpenStudy (anonymous):

VS=5.0V, VOH=4.5V, VIH=4.0V, VIL=1.5V, VOL=1.0V. What is the low noise margin (in Volts)? VIL-VOL=0.5 What is the high noise margin (in Volts)? VOH-VIH=0.5 What is the width of the forbidden region (in Volts)? VIH-VIL=2.5 Vout=Vs*RON/(RON+RPULLUP). (for nor and not gate) Vout=Vs*2*RON/(2*RON+RPULLUP). Vout is lower than VOL find rpullup What is the maximum power (in Watts) consumed by the inverter? VS^2/(RON+RPULLUP) 2.6315 What is the maximum power (in Watts) consumed by the NAND? VS^2/(2RON+RPULLUP) What is the maximum power (in Watts) consumed by the NOR? VS^2/(RON+RPULLUP)

OpenStudy (anonymous):

The static discipline required for this family is: VS=5.0V, VOH=4.5V, VIH=4.0V, VIL=1.5V, VOL=1.0V. What is the low noise margin (in Volts)? correct What is the high noise margin (in Volts)? correct What is the width of the forbidden region (in Volts)? correct Suppose that the threshold voltage for the MOSFET is VT=2.0V and RON=10000.0Ω. What is the minimum value of the pullup resistor RPuI (in Ohms) for which this inverter can obey the required static discipline? incorrect Now, consider the NAND gate of this family. What is the minimum value of the pullup resistor RPuA (in Ohms) for which this inverter can obey the required static discipline? incorrect How about the NOR gate of this family. What is the minimum value of the pullup resistor RPuO (in Ohms) for which this inverter can obey the required static discipline? incorrect Assume that we implemented this family with the minimum pullup resistors that you have already calculated. What is the maximum power (in Watts) consumed by the inverter? incorrect What is the maximum power (in Watts) consumed by the NAND? incorrect What is the maximum power (in Watts) consumed by the NOR?

OpenStudy (anonymous):

@lucky12345 nor gate power is wrong

OpenStudy (anonymous):

What is the minimum value of the pullup resistor RPuI (in Ohms) for which this inverter can obey the required static discipline? Answer: Rpull = [(Vs*Ron)/(VoL)] - Ron Now, consider the NAND gate of this family. What is the minimum value of the pullup resistor RPuA (in Ohms) for which this inverter can obey the required static discipline? Answer: Rpull = [(Vs*2*Ron)/(VoL)] - (2*Ron) How about the NOR gate of this family. What is the minimum value of the pullup resistor RPuO (in Ohms) for which this inverter can obey the required static discipline? Answer: Rpull = [(Vs*Ron)/(VoL)] - Ron What is the maximum power (in Watts) consumed by the inverter? Answer : Power = (Vs^2)/ (Rpull + Ron total) What is the maximum power (in Watts) consumed by the NAND? Answer : Power = (Vs^2)/ (Rpull + 2*Ron total) What is the maximum power (in Watts) consumed by the NOR? Answer : Power = (Vs^2)/ (Rpull + Ron||Ron)

OpenStudy (anonymous):

heja What is the minimum value of the pullup resistor RPuI (in Ohms) for which this inverter can obey the required static discipline? 40000 Now, consider the NAND gate of this family. What is the minimum value of the pullup resistor RPuA (in Ohms) for which this inverter can obey the required static discipline? 80000 How about the NOR gate of this family. What is the minimum value of the pullup resistor RPuO (in Ohms) for which this inverter can obey the required static discipline? 40000 What is the maximum power (in Watts) consumed by the inverter? 0.0005 What is the maximum power (in Watts) consumed by the NAND? 4.166666667e-4 What is the maximum power (in Watts) consumed by the NOR? 5.555555556e-4

OpenStudy (anonymous):

i am sorry power for the last part is Power = (Vs^2)/ (Rpull + Ron||Ron)

OpenStudy (anonymous):

im closing the question done it .... rap up brothas

OpenStudy (anonymous):

@lucky12345 no prob moving to question 3 i have got two wrongs nad no chance of error

OpenStudy (anonymous):

what do you mean by ron total

OpenStudy (anonymous):

Ron=10000 in ur case

OpenStudy (anonymous):

By Ron total i mean , that see how many mosfet are connected ? for e.g in nand , 2 mosfet are connected which means that Ron total = Ron + Ron

OpenStudy (anonymous):

@lucky12345 4.166666667e-4 is wrong for some reason?